# Analyzing Probabilistic Logic Shields for Multi-Agent Reinforcement Learning

### v2

@@ 4.3 Social Dilemmas in Stochastic, Mixed-Motive Environments (Extended Public Goods Game)
We use EPGG to investigate how PLS can promote cooperation in environments with stochastic social incentives. Our experiments address two central questions: ($i$) can agents learn to adopt normative strategies that reflect long-term expectations rather than short-term payoffs? and ($ii$) how does the effectiveness of shielding scale when only a subset of agents is shielded, especially under different degrees of parameter sharing? In the first experiment, we assess how shielding can steer agents toward the expected Nash equilibrium of the EPGG. In the second, we evaluate how the number of shielded agents and the degree of parameter sharing affect emergent cooperation and reward across a 5-agent population.
@@ 4.3.1 Learning Long-Term Normative Behavior
We begin with a two-player setting to isolate the effect of shielding on equilibrium selection under payoff uncertainty. This controlled setup enables a clear comparison between shielded agents following the expected Nash equilibrium and unshielded agents reacting to stochastic payoffs. Limiting the population removes confounding factors such as social influence, and allows for straightforward game-theoretic analysis, providing a clean baseline before scaling to larger populations.
@@
Figure 4 reports results for $\mu\in\{0.5,1,1.5,2.5,5.0\}$. Unshielded IPPO agents reliably converge to the instantaneous equilibrium – defecting when $f_{t}<1$ and cooperating when $f_{t}>2$. In contrast, SIPPO agents quickly converge to the desired behavior, leading to more prosocial behavior under uncertainty.
We observe that SIPPO agents exhibit lower reward variance and more stable cooperation compared to the baseline, especially in mixed-motive cases ($21<\mu<$ 2). For $\mu=0.5$, the shielded agents correctly learn to defect, but earn slightly less due to foregoing occasional cooperative gains from high $f_{t}$. Overall, shielded agents match or exceed the baseline in reward and demonstrate more consistent cooperation.
@@ 4.3.2 Partial Shielding and Parameter Sharing
Figure 3 shows grouped bar charts for $\mu\in\{0.5,1,1.5,2.5,5,7.5\}$. Results show a clear trend: increasing the number of shielded agents leads to higher overall cooperation and greater episodic reward in the mixed-motive range ($51<\mu<$ 5). Furthermore, this effect is amplified by parameter sharing: SACSPPO consistently achieves the same or higher cooperation than SCSPPO and SIPPO at the same shielded ratio. In fully cooperative settings ($\mu=7.5$), all configurations converge to high cooperation, but SACSPPO still exhibits the lowest variance. In competitive settings ($\mu=0.5$), cooperation remains low for the baseline algorithms, as expected, increasing with the number of shielded agents – this results in an increasingly lower mean reward with the benefit of the agents instead cooperating. Notably, in the competitive regime, SACSPPO exhibits the highest level of cooperation among the methods, indicating that shared learning with shielded agents can drive prosocial behavior in unshielded agents – even when it reduces their individual rewards.
@@ 4.4 Spatio-Temporal Coordination with Imperfect Information (Markov Stag-Hunt)
We use the Markov Stag-Hunt environment to investigate how PLS in SMARL shapes coordination in sequential, stochastic, and partially observable multi-agent settings. This environment extends the normal-form Stag-Hunt to a spatial domain where agents must learn when and how to cooperate over time [24]. Three experiments are conducted: In Section 4.4.1, we first study the effect of shield constraint strength on agent behavior. Then, in Section 4.4.2, we evaluate how cooperation scales with population size across a range of MARL algorithms. Finally, Section 4.4.3 explores how cooperation emerges when only a subset of agents is shielded, highlighting the role of parameter sharing in diffusing safety signals across the population.
@@ 4.4.1 Variation in Shield Strength
IPPO 18.74$\pm$4.98 0.09$\pm$0.11 5.50$\pm$2.04 13.71$\pm$4.97 16.07$\pm$5.80
SIPPO ($\mathcal{T}_{weak}$) 18.73$\pm$4.26 12.80$\pm$7.41 14.25$\pm$3.57 68.48$\pm$35.19 79.80$\pm$38.33 0.96$\pm$0.01
SIPPO ($\mathcal{T}_{strong}$) 11.33$\pm$4.08 77.32$\pm$8.82 11.07$\pm$3.25 386.86$\pm$46.50 393.07$\pm$65.27 0.95$\pm$0.01
To isolate the effect of constraint strength on emergent cooperation, we compare two shields: $\mathcal{T}_{weak}$ and $\mathcal{T}_{strong}$. These were constructed such that $\mathcal{T}_{strong}\subset\mathcal{T}_{weak}$ in terms of their safety constraints. The former constrains agents to cooperate only when both are near a stag, while the latter more strictly promotes cooperation regardless of position. Both shields are applied in a 2-agent setting using the SIPPO algorithm, with unshielded IPPO as a baseline.
In this experiment, analyzing reward or safety alone does not fully capture cooperative behavior, so the key metrics displayed in Table 3 are: total plant harvests per episode, successful stag hunts with another agent, and unsuccessful attempts to hunt a stag alone. Additionally, we show the total return for each episode, and mean safety for shielded agents. As expected, IPPO agents avoid cooperation (thereby avoiding risk) and consistently harvest plants for a small but reliable reward. SIPPO agents using $\mathcal{T}_{weak}$ show moderate improvement, achieving occasional stag hunts while maintaining similar plant collection. Their increased cooperation also results in more penalties, as agents are willing to risk solo hunts.
Under $\mathcal{T}_{strong}$, cooperation improves dramatically: agents prioritize stag hunting and harvest fewer plants, yet earn more than five times the episodic reward compared to SIPPO with $\mathcal{T}_{weak}$. This suggests that stronger shields not only induce safer behavior but also lead to better coordination and long-term returns.
@@ 4.4.2 Scaling Cooperation with Population Size
We next examine how cooperation and safety scale with group size when all agents are shielded using $\mathcal{T}_{strong}$. We vary the number of agents $n\in\{2,3,4,5\}$ and compare six algorithms: three unshielded baselines (IPPO, CSPPO, ACSPPO) and three shielded variants (SIPPO, SCSPPO, SACSPPO). Metrics include total plants harvested, stags hunted, solo hunt penalties, mean episodic safety, and per-agent reward.
Figure 5 shows that unshielded agents (IPPO, CSPPO, ACSPPO) consistently avoid cooperative behavior, harvesting plants for low-risk, low-reward outcomes. Notably, ACSPPO – despite its fully shared architecture – performs worst overall, with near-zero safety and minimal rewards, underscoring that parameter sharing alone is insufficient to induce prosocial behavior.
@@
These results demonstrate that PLS scales robustly with population size, especially when paired with full parameter sharing. SACSPPO leverages shared gradients and consistent safety signals to generalize cooperation more effectively across agents, even in more complex multi-agent settings.
@@ 4.4.3 Partial Shielding and Parameter Sharing
Finally, we investigate how cooperation emerges in partially shielded populations. Fixing the population size at $n=5$, we vary the number of shielded agents $k\in\{0,1,\dots,5\}$ and compare SIPPO, SCSPPO, and SACSPPO. All shielded agents use the strong shield $\mathcal{T}_{strong}$, and unshielded agents remain unconstrained. This experiment tests how shielded behavior propagates through shared parameters in populations where only a subset of agents receive explicit safety constraints.
@@
Together, these results demonstrate that PLS can induce prosocial coordination even under partial coverage, and that shared parameters are a key conduit for propagating safety-aligned behavior to unshielded agents.
@@ 5 Related Work
In the domain of safe (MA)RL, a few but significant strides have been made recently in developing frameworks that ensure safety during learning and execution [15]. Gu et al. [14], propose constrained Markov games and a constrained multi-agent policy gradient optimization algorithm (MACPO), however without the use of any formal language semantics. [35] use Mealy machines to approximate a world model to ensure safety, leading to some inherent symbolic interpretability. Other related concepts include dynamic shielding, for instance, ElSayed-Aly et al. [12] who introduce two frameworks that dynamically enforce safety constraints – our own method is similar to their factored shields. Banerjee et al. [3], Zhang et al. [40] and [38] are recent works that use the notion of (dynamic) model predictive shielding, which use ‘backup policies’ when an unsafe action is attempted by an agent. Though these are on par with SOTA results, these also do not provide the interpretability benefits of formal language semantics. Melcer et al. [20] explore decentralized shielding (similar to factored shields), i.e. allowing each agent to have its own shield, reducing computational overhead and enhancing scalability. This and ElSayed-Aly et al. [12] use LTL as formal language, though with a high knowledge requirement of the underlying MDP – this is in contrast to our use of a PL framework which requires only safety-related parts of the states. Notably, Subramanian et al. [29] integrate PL neural networks as the function approximator within agents to affect interpretable top-down control; however, our method differs significantly as the agents may use any function approximator, maximizing potential learning flexibility.
@@ 6 Discussion and Conclusion
    F: Training/reward curves.
@@ Appendix A Proofs
Let $\mathcal{T}$ be a ProbLog shield that, at a given state $s$, receives action facts action($a$) whose probabilities are the policy $\pi(a\mid s)$, and sensor facts $\ell$) that are deterministic in $s$ (i.e. we have perfect sensor information), i.e., $P_{\mathcal{T}}\!\bigl{(}\texttt{sensor}(\ell)=\mathrm{true}\bigr{)}\in\{0,1\}$ and $\texttt{sensor}(\ell)=\mathrm{true}\iff\ell\ \text{actually holds in }s.$
@@
[…]
@@
[…]
@@
Condition the ground program on action($a$). All remaining random facts are the deterministic $\ell$) atoms, so the program is now fully deterministic. Hence the queried probability is $1$ if and only if the rule for unsafe_next fails.
@@
Thus the conditional probability is always 00 or $1$, and partitioning actions by the value $1$ yields the stated set $A_{\mathrm{safe}}(s)$. ∎
@@
Because $S$, $A$ and therefore the post-shield state–action space $\{(s,a)\mid a\in A_{\text{safe}}(s)\}$ are finite, rewards are bounded, and the stepsizes $\{\eta_{t}\}$ satisfy $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$ (assumption iii), the classical stochastic-approximation proofs of Watkins and Dayan [36] for Q-learning or Sutton and Barto [30] for SARSA apply verbatim. A minor notational change is that the maximisation (in the off-policy case) is taken only over safe actions, but the contraction argument is identical. Under the usual further condition that every safe state–action pair is visited infinitely often (assumption i) the parameter sequence $\{\theta_{t}\}$ converges with probability 1 to the unique fixed point $Q^{\pi^{+}}$ of the Bellman operator corresponding to $\pi^{+}$.
@@
Note that the shield is state independent. Using the semantics of PL, we see that $P(\texttt{safe}|s,a^{\prime})=0$ for $a^{\prime}\neq a_{0}$, i.e. there is no state in which the safety of any other action apart from $a_{0}$ at all safe. Now, we can compute $\pi^{+}(a_{0}|s)$ and determine that $P(\texttt{safe}|s)=1-\sum\vec{a}_{-a_{0}}=a_{0}=\pi(a_{0}|s)$, leaving us with $\pi^{+}(a_{0}|s)=P_{\pi}(\texttt{safe}|s,a_{0})=1$. Thus, we get a safe deterministic policy (pure strategy) $\pi^{+}(s)=\langle 1,0,...\rangle$, sampling from which, the agent will always play $a_{0}$.∎
@@ B.1 Repeated NFGs (Stag-Hunt)
In these experiments both agents choose an action simultaneously and receive a reward determined by the payoff matrix. With reference to the generic payoff matrix in 7, any game with a payoff matrix with the following structure of payoffs is considered an instance of the Stag-Hunt: $a=b>d=e\geq g=h>c=f$. Each episode is a series of 25 repeated games. Thus the agents must learn to maximize the discounted reward attained over the whole episode [2]. The three Nash equilibria for the Stag-Hunt game instance as in Figure 2 are $(Stag,Stag)$, $(Hare,Hare)$ and a mixed Nash equilibrium where each agent plays Stag 60% of the time and Hare 40% of the time.
Figure 7: Generic payoff matrix
@@ B.4.1 Environment description
The adapted Markov Stag-Hunt environment used within this paper can be described in general as such: $n$ agents are placed in a square grid world of size $g\times g$. The agents have the following action set at each time step to move through the grid with respect to a global direction system: {UP, DOWN, LEFT, RIGHT, NONE}. There are $n_{stag}$ stags placed in the environment, as well as $n_{plant}$ plants, each occupying a single position each on the lattice. An exemplary visualisation of this environment can be seen in Figure 2. The agents can have either local or global observations, with or without a ‘self’ token, either as a 3D tensor or flattened one-hot encoded vector depending on the needs and configuration of the experiment.

### v3

@@ 4.3 Social Dilemmas in Stochastic, Mixed-Motive Environments (Extended Public Goods Game)
We use EPGG to investigate how PLS can promote cooperation in environments with stochastic social incentives. Our experiments address two central questions: ($i$) can agents be made to learn strategies that reflect long-term expectations rather than short-term payoffs?, and ($ii$) how does the effectiveness of shielding scale when only a subset of agents is shielded, and under different degrees of parameter sharing? In the first experiment, we assess how shielding can steer agents toward the expected Nash equilibrium of the EPGG. In the second, we evaluate how the degree of parameter sharing affects emergent cooperation.
@@ 4.3.1 Learning Long-Term Normative Behavior
We begin with a 2-player setting to isolate the effect of shielding on equilibrium selection under payoff uncertainty. This controlled setup enables a clear comparison between shielded agents learning long-term behavior versus unshielded agents reacting to stochastic payoffs. Limiting the population size removes confounding factors such as social influence and allows for direct game-theoretic analysis, providing a clean baseline before scaling to larger populations.
@@
Figure 4 reports results for $\mu\in\{0.5,1,1.5,2.5,5.0\}$. Unshielded IPPO agents reliably converge to the instantaneous equilibrium – defecting when $f_{t}<1$ and cooperating when $f_{t}>2$. In contrast, SIPPO agents quickly converge to the desired behavior, leading to more prosocial behavior under uncertainty. We observe that SIPPO agents exhibit lower reward variance and more stable cooperation compared to the baseline, especially in mixed-motive cases ($1<\mu<2$). For $\mu=0.5$, the shielded agents correctly learn to defect, but earn slightly less due to foregoing occasional cooperative gains from high $f_{t}$. Overall, shielded agents match or exceed the baseline in reward and demonstrate more consistent cooperation.
@@ 4.3.2 Partial Shielding and Parameter Sharing
Figure 3 shows grouped bar charts for $\mu\in\{0.5,1,1.5,2.5,5,7.5\}$. Results show a clear trend: increasing the number of shielded agents leads to higher overall cooperation and greater episodic reward in the mixed-motive range ($1<\mu<5$). Furthermore, this effect is amplified by parameter sharing: SACSPPO consistently achieves the same or higher cooperation than SCSPPO and SIPPO at the same shielded ratio. In fully cooperative settings ($\mu=7.5$), all configurations converge to high cooperation, but SACSPPO still exhibits the lowest variance. In competitive settings ($\mu=0.5$), cooperation remains low for the baseline algorithms, as expected, increasing with the number of shielded agents – this results in an increasingly lower mean reward with the benefit of the agents instead cooperating. Notably, in the competitive regime, SACSPPO exhibits the highest level of cooperation among the methods, indicating that shared learning with shielded agents can drive prosocial behavior in unshielded agents – even when it reduces their individual rewards.
@@ 4.4 Spatio-Temporal Coordination with Imperfect Information (Markov Stag-Hunt)
We use the Markov Stag-Hunt environment [30] to investigate how PLS-SMARL enables coordination in sequential, stochastic, and partially observable multi-agent settings. This environment extends the normal-form Stag-Hunt to a spatial domain where agents must learn when and how to cooperate. Three experiments are conducted: ($i$) we study the effect of shield constraint strength on agent behavior, ($ii$) we evaluate how cooperation scales with population size, (iii) finally we explore how cooperation emerges when only a subset of agents is shielded, highlighting the benefits of shielded parameter sharing.
@@ 4.4.1 Variation in Shield Strength
[…]
To isolate the effect of constraint strength on emergent cooperation, we compare two shields: $\mathcal{T}_{weak}$ and $\mathcal{T}_{strong}$. These were constructed such that $\mathcal{T}_{weak}\subset\mathcal{T}_{strong}$ in terms of their safety constraints. The former constrains agents to cooperate only when both are near a stag, while the latter more strictly promotes cooperation regardless of global position. Both shields are applied in a 2-agent setting using the SIPPO algorithm, with unshielded IPPO as a baseline.
In this experiment, analyzing just reward or safety does not fully capture cooperative behavior, so the key metrics displayed in Table 3 are: total plant harvests per episode, successful cooperative stag hunts, and unsuccessful attempts to hunt a stag alone. Additionally, we show the total return for each episode, and mean safety for shielded agents.
As expected, IPPO agents avoid cooperation (thereby avoiding risk) and steadily harvest plants for a small, reliable reward. SIPPO agents using $\mathcal{T}_{weak}$ show moderate improvement, achieving occasional stag hunts while maintaining similar plant collection. Their increased cooperation also results in more penalties, as agents are willing to risk solo hunts. Under $\mathcal{T}_{strong}$, cooperation improves dramatically: agents prioritize stag hunting and harvest fewer plants, yet earn more than five times the episodic reward compared to SIPPO with $\mathcal{T}_{weak}$. This suggests that well-designed stronger shields not only induce safer behavior, but may lead to better coordination and long-term returns.
@@ 4.4.2 Scaling Cooperation with Population Size
We next examine how cooperation and safety scale with group size when all agents are shielded using $\mathcal{T}_{strong}$. We vary the number of agents $n\in\{2,3,4,5\}$ and compare three unshielded baselines (IPPO, CSPPO, ACSPPO) and three shielded variants (SIPPO, SCSPPO, SACSPPO). Metrics include total plants harvested, stags hunted, solo hunt penalties, mean episodic safety, and per-agent reward.
Figure 5 shows that unshielded agents (IPPO, CSPPO, ACSPPO) consistently avoid cooperation, harvesting plants for low-risk, low-reward outcomes. Notably, ACSPPO – despite its fully shared architecture – performs worst overall, with near-zero safety and minimal rewards, underscoring that parameter sharing alone is insufficient to induce prosocial behavior.
@@
These results demonstrate that PLS scales robustly with population size. Notably, with full parameter sharing, SACSPPO leverages shared gradients and consistent safety signals to generalize cooperation more effectively across agents, even in complex multi-agent settings.
@@ 4.4.3 Partial Shielding and Parameter Sharing
Finally, we investigate cooperation in partially shielded populations. Fixing the population size at $n=5$, we vary the number of shielded agents $k\in\{0,1,\dots,5\}$ and compare SIPPO, SCSPPO, and SACSPPO. All shielded agents use the strong shield $\mathcal{T}_{strong}$, and unshielded agents remain unconstrained. This experiment tests how shielded behavior propagates through shared parameters in populations where only a subset of agents receive explicit safety constraints.
@@
Together, these results demonstrate that PLS can induce prosociality even under partial shielding, and that shared parameters are a key conduit for propagating safety-aligned behavior to unshielded agents.
@@ 5 Related Work
In safe (MA)RL, several recent frameworks address safety during learning and execution [17]. Gu et al. [16], propose constrained Markov games and MACPO, but without formal language semantics. Subramanian et al. [35] integrate PL neural networks as the function approximator within agents to affect interpretable top-down control; however, in our method, the agents may use any function approximator, maximizing potential learning flexibility. Waga et al. [41] use Mealy machines to approximate a world model to ensure safety, leading to some inherent symbolic interpretability. Other related concepts include dynamic shielding, for instance, ElSayed-Aly et al. [13] who introduce two frameworks that dynamically enforce safety constraints – our own method is similar to their factored shields. Banerjee et al. [4], Zhang et al. [47] and Xiao et al. [45] are recent works that use the notion of (dynamic) model predictive shielding. Though these show promising results, these do not benefit from the interpretability of formal language semantics. Melcer et al. [26] explore decentralized shielding (similar to factored shields), i.e. allowing each agent to have its own shield, reducing computational overhead and enhancing scalability. [26] and [13] use LTL as formal language, though with a high knowledge requirement of the underlying MDP – this is in contrast to our use of PLS which requires only safety-related parts of the states. Finally, some methods, such as Wang et al. [42], enforce safety through natural language constraints; in contrast, our approach builds on PLS to provide stronger formal guarantees.
@@ 6 Discussion and Conclusion
{ack}
This publication is part of the project ‘Hybrid Intelligence: augmenting human intellect’ (https://hybrid-intelligence-centre.nl,) with project number 024.004.022 of the research programme ‘Gravitation’ which is (partly) financed by the Dutch Research Council (NWO).
@@
    F: Partial training/reward curves.
@@ Appendix A Proofs
Let $\mathcal{T}$ be a ProbLog shield that, at a given state $s$, receives action facts action($a$) whose probabilities are the policy $\pi(a\mid s)$, and sensor facts sensor($\ell$) that are deterministic in $s$ (i.e. we have perfect sensor information), i.e., $P_{\mathcal{T}}\!\bigl{(}\texttt{sensor}(\ell)=\mathrm{true}\bigr{)}\in\{0,1\}$ and $\texttt{sensor}(\ell)=\mathrm{true}\iff\ell\ \text{actually holds in }s.$
@@
$\texttt{unsafe\_next}\;:\!-\;\texttt{action}(a),\,\texttt{sensor}(\ell_{a}).$ (4)
@@
$\texttt{safe\_next}\;:\!-\;\lnot\texttt{unsafe\_next},$
@@
Condition the ground program on action($a$). All remaining random facts are the deterministic sensor($\ell$) atoms, so the program is now fully deterministic. Hence the queried probability is $1$ if and only if the rule for unsafe_next fails.
@@
Thus the conditional probability is always 0 or $1$, and partitioning actions by the value $1$ yields the stated set $A_{\mathrm{safe}}(s)$. ∎
@@
Because $S$, AA and therefore the post-shield state–action space $\{(s,a)\mid a\in A_{\text{safe}}(s)\}$ are finite, rewards are bounded, and the stepsizes $\{\eta_{t}\}$ satisfy $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$ (assumption iii), the classical stochastic-approximation proofs of Watkins and Dayan [43] for Q-learning or Sutton and Barto [36] for SARSA apply verbatim. A minor notational change is that the maximisation (in the off-policy case) is taken only over safe actions, but the contraction argument is identical. Under the usual further condition that every safe state–action pair is visited infinitely often (assumption i) the parameter sequence $\{\theta_{t}\}$ converges with probability 1 to the unique fixed point $Q^{\pi^{+}}$ of the Bellman operator corresponding to $\pi^{+}$.
@@
Note that the shield is state independent. Using the semantics of PL, we see that $P(\texttt{safe}|s,a^{\prime})=0$ for $a^{\prime}\neq a_{0}$, i.e. there is no state in which the safety of any other action apart from $a_{0}$ at all safe. Now, we can compute $\pi^{+}(a_{0}|s)$ and determine that $P(\texttt{safe}|s)=1-\sum\vec{a}_{-a_{0}}=a_{0}=\pi(a_{0}|s)$, leaving us with $\pi^{+}(a_{0}|s)=P_{\pi}(\texttt{safe}|s,a_{0})=1$. Thus, we get a safe deterministic policy (pure strategy) $\pi^{+}(s)=\langle 1,0,$$\rangle$, sampling from which, the agent will always play $a_{0}$.∎
@@ B.1 Repeated NFGs (Stag-Hunt)
In these experiments both agents choose an action simultaneously and receive a reward determined by the payoff matrix. With reference to the generic payoff matrix in 7, any game with a payoff matrix with the following structure of payoffs is considered an instance of the Stag-Hunt: $a=b>d=e\geq g=h>c=f$. Each episode is a series of 25 repeated games. Thus the agents must learn to maximize the discounted reward attained over the whole episode [3]. The three Nash equilibria for the Stag-Hunt game instance as in Figure 2 (main paper) are $(Stag,Stag)$, $(Hare,Hare)$ and a mixed Nash equilibrium where each agent plays Stag 60% of the time and Hare 40% of the time.
@@ B.4.1 Environment description
The adapted Markov Stag-Hunt environment used within this paper can be described in general as such: $n$ agents are placed in a square grid world of size $g\times g$. The agents have the following action set at each time step to move through the grid with respect to a global direction system: {UP, DOWN, LEFT, RIGHT, NONE}. There are $n_{stag}$ stags placed in the environment, as well as $n_{plant}$ plants, each occupying a single position each on the lattice. The agents can have either local or global observations, with or without a ‘self’ token, either as a 3D tensor or flattened one-hot encoded vector depending on the needs and configuration of the experiment.